# Understanding Transformers via $N$-gram Statistics

### v1

@@ Appendix A Choice of Distance Measure
We choose variational distance since it is a symmetric and bounded distance function (unlike the KL divergence). Symmetry means we do not have to make a choice between computing the distance between model predictions and rule predictions or vice versa. Boundedness ensures that when we measure average distance across tokens, large outliers do not dominate the average. In fact, for the KL divergence, since $KL(p||q)$ is infinite when $p>0$ wherever $q=0$, were we to use KL divergence, we would have to set $p$ equal to rule predictions and $q$ equal to model predictions (since rule predictions are typically sparse). To avoid such constraints and potential pathologies, we choose the variational distance as our metric. We find that other measures like Jensen-Shannon distance give similar results. It is worth noting that while the $L^{\infty}$-metric often gives similar results, it has a failure mode when comparing two very high entropy distributions. If $p$ and $q$ are two distributions such that $p_{i}$ and $q_{i}$ are all small, then their $L^{\infty}$ distance will be small even though their variational distance can be large.
@@ Appendix B Additional Experimental Details
Table 3: Model Specifications
We use a linear learning rate warmup of 1000 steps up to a maximum value of $2\times 10^{-4}$ and then use a cosine learning rate decay. We use weighted Adam optimizer (Loshchilov and Hutter,, 2017) with weight decay $10^{-4}$. Our models are trained using TPU accelerators. The 160M models use 16 TPU accelerators while the 1.4B models use 64 TPU accelerators (to exploit data parallelism) per run. We use a batch size of 128 sequences with each sequence consisting of 2048 tokens.
@@
For TinyStories experiments, we train 160M models for 4 epochs except for the overfitting experiments where we train 1.4B models for 10 epochs. We use the train and validation splits provided by HuggingFace [^15] Available at https://huggingface.co/datasets/roneneldan/TinyStories. For Wikipedia experiments, we train a 1.4B model for a single epoch. We have train and validation splits based on using choosing random sets of disjoint documents. Our Wikipedia train set has 4.4B tokens. In places where we perform several training runs (Section 5), the only source of variance (randomness) among the runs are different dataset shuffles.
@@ B.1 $N$-gram statistics
Appendix C Additional Approximation-Variance Association Analysis
@@ C.1 Full-context vs Subcontext
As noted in footnote 11, there is usually a mismatch between the contexts that $N$-gram rules and LLMs receive during training: the latter can receive very long contexts (up to one less than the number of tokens in a document) while the former typically receives very short contexts (in our case, up to $7$ tokens). Concretely, while a bigram model is trained on consecutive pairs of tokens $(c,t)$, an LLM is rarely trained so as to optimize $p(t|c)$. Indeed, given a training sequence $x$, only the target for the first token of $x$ has context consisting of a single token; the other targets will have more tokens of context accordingly. Thus, it is unclear how well LLM predictions $p(t|c)$ should match bigram rule predictions as $c$ varies over the vocabulary set, since LLMs almost always receive $c$ within a much larger context. More generally, it is unclear how well $p(t|C)$ matches $p_{\textrm{full}}(t|C)$. Nevertheless, because in practice LLMs learn how to use context effectively, LLMs manage to learn $p(t|C)$ despite being optimized for $p(t|\tilde{C})$ with $\tilde{C}$ a context containing $C$ as a suffix.
@@
Figure 6: Comparison with TinyStories bigram model. We evaluate transformer models (trained with either full context or context length equal to one) on all 22.8K distinct unigrams of TinyStories and record the corresponding variational distance with the bigram rule. Grouping unigrams based on count and averaging the variational distances result in the above scatterplots.
In both cases, we have the same pattern of increased count leads to lower rule distance. However, the context length equal to one transformer has much lower distances since it cannot learn anything else other than the bigram rule. The difference between the variational differences of the two models is thus a measure of the dilution an LLM has in learning a bigram rule owing to receiving surrounding context.
@@ C.2 TinyStories Unigram Context
We repeat Figure 2 for the simplest case of unigram context. In this case, there is only one rule (the bigram rule) and so there are only three plots to consider. It turns out also the correlation between optimal distance and count is slightly stronger than with model variance. However, given the unigram context case is extreme (in the sense that there is only a single token of context), we treat this case as an edge case.
slope $=-0.07$ ; $R^{2}=0.32$
@@
slope = $-0.13$, $R^{2}=0.64$
Figure 7: TinyStories $1$-grams. Every point in the above plots represents a $1$-gram context (all 22.8K from TinyStories). Shaded regions are plots obtained by bucketing along the x-axis and computing one standard deviation within the mean along the y-axis. Slope and $R^{2}$ values of plots are with respect to the linear fit of the data given by their axes. Optimal rule distances and model variances are computed with respect to five model runs. (a): model variance vs count of $C$ (b): $d(p(t|C),p_{\textrm{full}}(t|C))$ vs model variance (c): $d(p(t|C),p_{\textrm{full}}(t|C))$ vs count
@@ C.3 Tinystories Bigram Context
Figure 8: TinyStories full-context bigrams. Every point in the above plots represents a full-context bigram $C$ from among the 691 distinct ones in TinyStories. Points are colored by which $N$-gram rule is the optimal rule, among those in $\mathcal{R}^{\textrm{all}}_{2}$, for transformer prediction for the given context. Shaded regions are plots obtained by bucketing along the x-axis and computing one standard deviation within the mean along the y-axis. (a): $d(p(\cdot|C),p_{\textrm{trigram}}(\cdot|C))$ vs count of $C$. (b): $d(p(\cdot|C),p_{\textrm{trigram}}(\cdot|C))$ vs trigram-model predictions. (c): Optimal rule distance vs the greater of the bigram count of $C$ and the unigram count of $C_{-1}$. (d): Similar to upper right but now the y-axis is optimal rule distance. Five model runs are used to compute optimal rule distance and model variance.
@@ C.4 Wikipedia 6-gram contexts
We plot the analog of Figure 2 but for contexts consisting of $6$-grams from Wikipedia. We also subsample as before, from logarithmically spaced buckets, for a total of around 6.8K total contexts. We get nearly identical behavior as with TinyStories. Our approximation-variance association is thus not specific to small datasets like TinyStories.
slope $=-0.23$ ; $R^{2}=0.26$
@@
Figure 9: Wikipedia 6-grams. Every point in the above plots represents a $6$-gram context. Shaded regions are plots obtained by bucketing along the x-axis and computing one standard deviation within the mean along the y-axis. Slope and $R^{2}$ values of plots are with respect to the linear fit of the data given by their axes. Optimal rule distances and model variances are computed with respect to five model runs. (a): $d(p(t|C),p_{\textrm{full}}(t|C))$ vs count of $C$. (b): $d(p(t|C),p_{\textrm{full}}(t|C))$ vs model variance. (c): model variance vs count of $C$. (d): similar to (b) but now the y-axis is optimal rule distance of the optimal rule from $\mathcal{R}^{\textrm{suffix}}_{6}$.
@@ D.1 Rule Approximation: A Closer Look
We supplement Table 2 with Table 4 to show how optimal rule distances decrease with increasing rule strength. This is to preclude a trivial situation in which by having sufficiently many rules (say a one-hot distribution for every vocabulary token), one can have a ruleset that for any model prediction always returns an optimal rule with 100% top-1 accuracy! Such coarse rules will not in general yield small optimal distances however [^18] The variational distance between a one-hot distribution and a distribution which is uniform on $n$ tokens is at least $\frac{n-1}{n}$. Thus, whenever an LLM has at least two roughly valid options, we expect a one-hot distribution to be at least of distance $0.5$ from the LLM prediction. and our variational distances decreasing in Table 4 shows that our rulesets are truly better approximating the predictions with increasing strength.
ruleset / $M$ 1 2 3 4 5 6 7
$\mathcal{R}^{\textrm{all}}_{M}$ 0.738 0.596 0.507 0.433 0.369 0.315 0.273
$\mathcal{R}^{\textrm{subgram}}_{M}$ 0.738 0.597 0.512 0.448 0.398 0.361 0.334
$\mathcal{R}^{\textrm{suffix}}_{M}$ 0.738 0.598 0.519 0.464 0.425 0.399 0.381
@@
Table 4: Average Optimal Rule distance. We look at the average optimal rule distance with LLM predictions for rules of varying strength and maximum context length. We compute this average over each token prediction from 100 random TinyStories validation stories (around 22K tokens total).
D.2 Rule Approximation: Another Interpretation
Our rule approximation is formulated as a retrodictive procedure in which after an LLM prediction is made, one uses the optimization procedure defined in Section 5 to select an optimal rule for describing the prediction. This retrodictive viewpoint however can be replaced with an alternative interpretation:
Regard the LLM output next-token distribution as a value-vector and each rule prediction as a key-vector in probability space. We use nearest-neighbors with respect to variational distance to select the closest key to the given LLM value-vector. This key then becomes the resulting predicted probability distribution of this joint system of an LLM plus $N$-gram rules. The joint system is then a predictive model that forces LLM predictions through a “bottleneck layer" of a small set of $N$-gram rules. Our results from Section 7 can be interpreted as saying that such an $N$-gram bottleneck layer achieves 78% fidelity with respect to the original LLM predictions on TinyStories as measured by top-1 accuracy.
D.3 TinyStories

### v2

@@ Appendix A Choice of Distance Measure
We choose variational distance since it is a symmetric and bounded distance function (unlike the KL divergence). Symmetry means we do not have to make a choice between computing the distance between model predictions and rule predictions or vice versa. Boundedness ensures that when we measure average distance across tokens, large outliers do not dominate the average. In fact, for the KL divergence, since $KL(p||q)$ is infinite when $p>0$ whenever $q=0$, were we to use KL divergence, we would have to set $p$ equal to rule predictions and $q$ equal to model predictions (since rule predictions are typically sparse). To avoid such constraints and potential pathologies, we choose the variational distance. Another possible choice is the Jensen-Shannon distance, but we found it gives similar results to variational distance.
It is worth noting that while the $L^{\infty}$ metric
$d_{L^{\infty}}(p,q)=\max_{\alpha}|p_{\alpha}-q_{\alpha}|$ (14)
often gives slightly better results for the rule-approximation analysis of Section 7, it has a failure mode when comparing two very high entropy distributions. If $p$ and $q$ are two distributions such that $p_{\alpha}$ and $q_{\alpha}$ are all small, then their $L^{\infty}$ distance will be small even though their variational distance can be large. Thus, the $L^{\infty}$ distance is not suitable for the curriculum learning analysis analysis in Section 6.1, since at initialization when the model makes uniform predictions on tokens, it will have low $L^{\infty}$ distance with many bad $N$-gram rules that are high entropy. Hence, for the sake of using a consistent metric in the main body of the paper, we use the variational distance when comparing probability distributions, although as the numbers of Section D show, for well-trained models the $L^{\infty}$ distance often yields better rule approximation.
@@ Appendix B Additional Experimental Details
We use a linear learning rate warmup of 1000 steps up to a maximum value of $2\times 10^{-4}$ and then use a cosine learning rate decay. We use weighted Adam optimizer (Loshchilov and Hutter,, 2017) with weight decay $10^{-4}$. Our models are trained using TPU accelerators. The 160M and 420M models use 16 TPU accelerators while the 1.4B models use 64 TPU accelerators per run. We use a batch size of 128 sequences with each sequence consisting of 2048 tokens.
@@
For TinyStories experiments, we train 160M models for 4 epochs except for the overfitting experiments in Section 6.2 where we train 1.4B models for 10 epochs. We use the train and validation splits provided by HuggingFace [^16] Available at https://huggingface.co/datasets/roneneldan/TinyStories. For Wikipedia experiments, we train a 1.4B model for a single epoch. We have train and validation splits based on using choosing random sets of disjoint documents. Our Wikipedia train set has 4.4B tokens. In places where we perform several training runs (Section 5), the only source of variance (randomness) among the runs are different dataset shuffles. The only exception to the above is in Section D.3, where all model sizes are trained for 1 epoch (for TinyStories, we observed overfitting of the 1.4B model around 4 epochs so we switch to 1 epoch for all model sizes for a fair comparison).
@@ B.1 $N$-gram statistics
Appendix C Additional Approximation Criterion Analyses
@@ C.1 Full-context vs Subcontext
As noted in footnote 12, there is usually a mismatch between the contexts that $N$-gram rules and LLMs receive during training: the latter can receive very long contexts (up to one less than the number of tokens in a document) while the former typically receives very short contexts (in our case, up to $7$ tokens). Concretely, while a bigram model is trained on consecutive pairs of tokens $(c,t)$, an LLM is rarely trained so as to optimize $p(t|c)$. Indeed, given a training sequence $x$, only the target for the first token of $x$ has context consisting of a single token; the other targets will have more tokens of context accordingly. Thus, it is unclear how well LLM predictions $p(t|c)$ should match bigram rule predictions as $c$ varies over the vocabulary set, since LLMs almost always receive $c$ within a much larger context. More generally, it is unclear how well $p(t|C)$ matches $p_{\textrm{full}}(t|C)$. Nevertheless, because in practice LLMs learn how to use context effectively, LLMs manage to learn $p(t|C)$ despite being optimized for $p(t|\tilde{C})$ with $\tilde{C}$ a context containing $C$ as a suffix.
@@
Figure 6: Comparison with TinyStories bigram model. We evaluate transformer models (trained with either full context or context length equal to one) on all 22.8K distinct unigrams of TinyStories and record the corresponding variational distance with the bigram rule. Grouping unigrams based on count and averaging the variational distances result in the above scatterplots. Model size: 160M.
In both cases, we have the same pattern of increased count leads to lower rule distance. However, the transformer trained with context length equal to one has much lower distances since it cannot learn anything else other than the bigram rule. The difference between the variational differences of the two models is thus a measure of the dilution an LLM has in learning a bigram rule owing to receiving surrounding context.
@@ C.2 TinyStories Unigram Context
We repeat Figure 2 for the simplest case of unigram context. In this case, there is only one rule (the bigram rule) and so there are only three plots to consider.
slope $=-0.02$ ; $R^{2}=0.26$
@@
slope = $-0.07$, $R^{2}=0.65$
Figure 7: TinyStories $1$-grams. Every point in the above plots represents a $1$-gram context (all 22.8K from TinyStories). Shaded regions are plots obtained by bucketing along the x-axis and computing one standard deviation within the mean along the y-axis. Slope and $R^{2}$ values of plots are with respect to the linear fit of the data given by their axes. Optimal rule distances and model variances are computed with respect to five model runs. (a): model variance vs count of $C$ (b): $d(p(t|C),p_{\textrm{full}}(t|C))$ vs model variance (c): $d(p(t|C),p_{\textrm{full}}(t|C))$ vs count. Model size: 160M.
@@ C.3 Tinystories Bigram Context
Figure 8: TinyStories full-context bigrams. Every point in the above plots represents a full-context bigram $C$ from among the 691 distinct ones in TinyStories. Points are colored by which $N$-gram rule is the optimal rule, among those in $\mathcal{R}^{\textrm{all}}_{2}$, for transformer prediction given $C$. Shaded regions are plots obtained by bucketing along the x-axis and computing one standard deviation within the mean along the y-axis. (a): $d(p(\cdot|C),p_{\textrm{trigram}}(\cdot|C))$ vs count of $C$. (b): $d(p(\cdot|C),p_{\textrm{trigram}}(\cdot|C))$ vs model variance. (c): Optimal rule distance vs the greater of the bigram count of $C$ and the unigram count of $C_{-1}$. (d): Similar to upper right but now the y-axis is optimal rule distance. Five model runs are used to compute optimal rule distance and model variance. Model size: 160M.
@@ C.4 Wikipedia 6-gram contexts
We plot the analog of Figure 2 in Figure 9 but for contexts consisting of $6$-grams from Wikipedia. We also subsample as before, from logarithmically spaced buckets, for a total of around 6.8K total contexts. We get nearly identical behavior as with TinyStories. Our Approximation Criterion is thus not specific to small datasets like TinyStories.
slope $=-0.06$ ; $R^{2}=0.45$
@@
Figure 9: Wikipedia 6-grams. Every point in the above plots represents a $6$-gram context. Shaded regions are plots obtained by bucketing along the x-axis and computing one standard deviation within the mean along the y-axis. Slope and $R^{2}$ values of plots are with respect to the linear fit of the data given by their axes. Optimal rule distances and model variances are computed with respect to five model runs. (a): $d(p(t|C),p_{\textrm{full}}(t|C))$ vs count of $C$. (b): $d(p(t|C),p_{\textrm{full}}(t|C))$ vs model variance. (c): model variance vs count of $C$. (d): similar to (b) but now the y-axis is optimal rule distance of the optimal rule from $\mathcal{R}^{\textrm{suffix}}_{6}$. Model size: 160M.
@@ D.1 TinyStories
We supplement Table 2 with Table 4 to show how optimal rule distances decrease with increasing rule strength. This is to preclude a trivial situation in which by having sufficiently many rules (say a one-hot distribution for every vocabulary token), one can have a ruleset that for any model prediction always returns an optimal rule with 100% top-1 accuracy! Such coarse rules will not in general yield small optimal distances however [^19] Both the variational and $L^{\infty}$ distances between a one-hot distribution and a distribution which is uniform on $n$ tokens are at least $\frac{n-1}{n}$. Thus, whenever an LLM has at least two roughly valid options, we expect a one-hot distribution to be at least of distance $0.5$ from the LLM prediction. and our variational distances decreasing in Table 4 shows that our rulesets are truly better approximating the predictions with increasing strength.
We also include the analog of Tables 2 and (4 but with the variational distance replaced with the $L^{\infty}$ distance in Tables 5 and 6. We see that in fact the top-1 accuracy numbers are slightly better in this case.
@@
Table 4: Optimal rule distance (TinyStories, variational distance). We look at the average optimal rule distance with LLM predictions for rules of varying strength and maximum context length $M$. We compute this average over each token prediction from 100 random TinyStories validation stories (around 22K tokens total). Model size: 160M.
ruleset / context length 1 2 3 4 5 6 7
$\mathcal{R}^{\textrm{all}}_{M}$ 0.738 0.597 0.507 0.434 0.37 0.316 0.274
$\mathcal{R}^{\textrm{subgram}}_{M}$ 0.738 0.598 0.513 0.449 0.399 0.362 0.335
$\mathcal{R}^{\textrm{suffix}}_{M}$ 0.738 0.598 0.519 0.465 0.425 0.399 0.381
$\mathrm{backoff}_{M}$ 0.738 0.606 0.539 0.503 0.482 0.472 0.466
Table 5: Top-1 accuracy of optimal rule (TinyStories, $L^{\infty}$ distance). The analogue of Table 2 but with $L^{\infty}$ distance instead of variational distance for selecting the optimal rule. Model size: 160M.
ruleset / context length 1 2 3 4 5 6 7
$\mathcal{R}^{\textrm{all}}_{M}$ 30.0 45.1 55.0 63.2 69.6 75.0 78.9
$\mathcal{R}^{\textrm{subgram}}_{M}$ 30.0 44.7 53.5 60.4 65.1 68.5 71.2
$\mathcal{R}^{\textrm{suffix}}_{M}$ 30.0 44.5 52.6 58.0 61.4 63.6 65.2
$\mathrm{backoff}_{M}$ 30.0 42.5 48.7 52.7 54.6 55.9 56.7
Table 6: Optimal rule distance (TinyStories, $L^{\infty}$ distance). The analogue of Table 4 but with $L^{\infty}$ distance instead of variational distance. Model size: 160M.
ruleset / context length 1 2 3 4 5 6 7
$\mathcal{R}^{\textrm{all}}_{M}$ 0.557 0.443 0.371 0.309 0.253 0.207 0.170
$\mathcal{R}^{\textrm{subgram}}_{M}$ 0.557 0.444 0.376 0.322 0.28 0.248 0.225
$\mathcal{R}^{\textrm{suffix}}_{M}$ 0.557 0.446 0.383 0.338 0.305 0.282 0.267
$\mathrm{backoff}_{M}$ 0.557 0.456 0.407 0.385 0.378 0.379 0.382
D.2 Wikipedia
We present analogous results for Section 7 using a 1.4B model trained on Wikipedia. In Table 7 we present the analogue of Table 2 on ten holdout Wikipedia chunks (a total of $10\times 2048$ tokens). The top-1 accuracy when using optimal rules from $\mathcal{R}^{\textrm{all}}_{7}$ and the $L^{\infty}$ distance for rule selection is 67.7%. As with TinyStories, we see significant gains in accuracy when we increase rule strength. Achieving the number 67.7% (versus the corresponding 78.9% number for TinyStories from Table 5), perhaps a surprisingly a high score, is the result of two competing factors: on the one-hand, Wikipedia is a more complex dataset (which makes prediction harder), while on the other hand, the training data has more $N$-grams and thus more rules. Our model achieves 54.9% top-1 accuracy on the 10 holdout Wikipedia chunks (Table 12) which is substantially lower than the top-1 accuracy of the optimal rule.
Table 7: Top-1 accuracy of optimal rule (Wikipedia, $L^{\infty}$ distance). We look at the average top-1 accuracy between optimal rule and LLM predictions for rules of varying strength and maximum context length. We compute this average over each token prediction from 10 holdout Wikipedia sequences each consisting of 2048 tokens. Model size: 1.4B.
ruleset / context length 1 2 3 4 5 6 7
$\mathcal{R}^{\textrm{all}}_{M}$ 24.3 39.7 49.8 55.5 60.3 64.3 67.7
$\mathcal{R}^{\textrm{subgram}}_{M}$ 24.3 39.2 48.3 52.7 55.5 57.6 59.0
$\mathcal{R}^{\textrm{suffix}}_{M}$ 24.3 38.8 47.1 50.3 51.5 52.3 52.7
$\textrm{backoff}_{M}$ 24.3 36.8 42.9 43.9 43.7 43.8 43.9
Table 8: Optimal rule distance (Wikipedia, $L^{\infty}$ distance). We look at the average $L^{\infty}$ distance between optimal rule and LLM predictions for rules of varying strength and maximum context length. We compute this average over each token prediction from 10 holdout Wikipedia sequences each consisting of 2048 tokens. Model size: 1.4B.
ruleset / context length 1 2 3 4 5 6 7
$\mathcal{R}^{\textrm{all}}_{M}$ 0.48 0.369 0.29 0.237 0.202 0.173 0.150
$\mathcal{R}^{\textrm{subgram}}_{M}$ 0.48 0.371 0.295 0.249 0.226 0.209 0.198
$\mathcal{R}^{\textrm{suffix}}_{M}$ 0.48 0.375 0.303 0.265 0.250 0.242 0.238
$\mathrm{backoff}_{M}$ 0.48 0.394 0.359 0.368 0.387 0.398 0.405
Table 9: Top-1 accuracy of optimal rule (Wikipedia, variational distance). We look at the average top-1 accuracy between optimal rule and LLM predictions for rules of varying strength and maximum context length. We compute this average over each token prediction from 10 holdout Wikipedia sequences each consisting of 2048 tokens. Model size: 1.4B.
ruleset / context length 1 2 3 4 5 6 7
$\mathcal{R}^{\textrm{all}}_{M}$ 24.3 39.2 48.8 54.4 58.5 61.9 65.0
$\mathcal{R}^{\textrm{subgram}}_{M}$ 24.3 39.0 47.9 52.3 54.9 56.8 58.3
$\mathcal{R}^{\textrm{suffix}}_{M}$ 24.3 38.7 46.9 50.3 51.5 52.4 52.8
$\textrm{backoff}_{M}$ 24.3 36.8 42.9 43.9 43.7 43.8 43.9
Table 10: Optimal rule distance (Wikipedia, variational distance). We look at the average variational distance between optimal rule and LLM predictions for rules of varying strength and maximum context length. We compute this average over each token prediction from 10 holdout Wikipedia sequences each consisting of 2048 tokens. Model size: 1.4B.
ruleset / context length 1 2 3 4 5 6 7
$\mathcal{R}^{\textrm{all}}_{M}$ 0.768 0.635 0.543 0.484 0.446 0.413 0.388
$\mathcal{R}^{\textrm{subgram}}_{M}$ 0.768 0.636 0.549 0.498 0.472 0.453 0.440
$\mathcal{R}^{\textrm{suffix}}_{M}$ 0.768 0.638 0.556 0.513 0.497 0.488 0.483
$\mathrm{backoff}_{M}$ 0.768 0.656 0.609 0.597 0.598 0.599 0.600
D.3 Model Scaling
In this section, we present results about how our rule approximation changes with model size. For both TinyStories and Wikipedia, we train models of size 160M, 420M, and 1.4B for one epoch [^20] Since the 1.4B starts showing signs of overfitting around 4 epochs on TinyStories, we train models for 1 epoch in this section unlike 4 epochs elsewhere..
Table 11: TinyStories metrics. How average cross entropy loss, top-1 accuracy, and model distance to the ground truth scale with model size on a holdout set of 100 stories.
model size eval loss eval acc eval distance
160M 1.43 63.2 0.485
420M 1.28 65.9 0.452
1.4B 1.22 66.9 0.439
Table 12: Wikipedia metrics. How average cross entropy loss, top-1 accuracy, and model distance to the ground truth scale with model size on a holdout set of 10 Wikipedia chunks.
model size eval loss eval acc eval distance
160M 2.63 50.0 0.617
420M 2.43 52.3 0.590
1.4B 2.26 54.9 0.562
In Tables 11 and 12, there is a clear trend towards improved model performance with scale: we obtain lower cross entropy loss, higher top-1 accuracy, and lower model distance to the ground truth distribution regarded as a one-hot distribution [^21] It would be more proper to aggregate statistics over the holdout set to compute a ground truth distribution that takes into the account the relative frequencies of the next token given the context. However, since most contexts will be unique, this more refined computation will not affect the corresponding result much.. Note that for the latter, distance from model predictions to the ground truth distribution is the same with respect to the variational or $L^{\infty}$ distance and is given by $1-p$ where $p$ is the probability assigned by the model to the ground truth.
On the other hand, entries in Tables 13-20 measuring the changes in rule approximation with scale are much more modest (i.e. are more stable) by comparison. They also show mixed results, with $L^{\infty}$ results worsening but some variational distance results slightly improving with scale (for large context length rules). We leave a more thorough investigation of how model scale affects approximability by $N$-gram rules to future work.
Table 13: Top-1 accuracy of optimal rule with model scale (TinyStories, variational distance). How top-1 accuracy of the optimal rule varies with model size and rule context length. Optimal rule from is selected from $\mathcal{R}^{\textrm{all}}_{7}$ using variational distance.
model size / context length 1 2 3 4 5 6 7
160M 31.4 47.1 56.7 64.3 69.8 74.5 77.9
420M 30.7 45.9 55.5 63.3 69.2 74.1 77.9
1.4B 30.5 45.9 55.3 63.4 69.6 74.3 78.2
Table 14: Optimal rule distance with model scale (TinyStories, variational distance). How optimal rule distance varies with model size and rule context length. Optimal rule is selected from $\mathcal{R}^{\textrm{all}}_{7}$ using variational distance.
model size / context length 1 2 3 4 5 6 7
160M 0.692 0.545 0.46 0.398 0.347 0.306 0.275
420M 0.711 0.566 0.479 0.411 0.355 0.310 0.274
1.4B 0.718 0.574 0.486 0.417 0.359 0.311 0.274
Table 15: Top-1 accuracy of optimal rule with model scale (TinyStories, $L^{\infty}$ distance). How top-1 accuracy of the optimal rule varies with model size and rule context length. Optimal rule is selected from $\mathcal{R}^{\textrm{all}}_{7}$ using $L^{\infty}$ distance.
model size / context length 1 2 3 4 5 6 7
160M 31.4 47.5 57.5 65.2 71.4 76.1 79.6
420M 30.7 46.3 56.3 64.2 70.5 75.3 79.2
1.4B 30.5 46.2 56.0 64.1 70.5 75.3 79.3
Table 16: Optimal rule distance with model scale (TinyStories, $L^{\infty}$ distance). How optimal rule distance varies with model size and rule context length. Optimal rule is selected from $\mathcal{R}^{\textrm{all}}_{7}$ using $L^{\infty}$ distance.
model size / context length 1 2 3 4 5 6 7
160M 0.482 0.368 0.301 0.248 0.204 0.169 0.141
420M 0.511 0.398 0.329 0.272 0.223 0.184 0.153
1.4B 0.522 0.408 0.338 0.280 0.230 0.189 0.157
Table 17: Top-1 accuracy of optimal rule with model scale (Wikipedia, variational distance). How top-1 accuracy of the optimal rule varies with model size and rule context length. Optimal rule is selected from $\mathcal{R}^{\textrm{all}}_{7}$ using variational distance.
model size / context length 1 2 3 4 5 6 7
160M 25.9 40.8 49.5 54.2 57.7 61.0 63.8
420M 24.7 39.8 49.0 54.2 58.1 61.6 64.3
1.4B 24.3 39.2 48.8 54.4 58.5 61.9 65.0
Table 18: Optimal rule distance with model scale (Wikipedia, variational distance). How optimal rule distance varies with model size and rule context length. Optimal rule is selected from $\mathcal{R}^{\textrm{all}}_{7}$ using variational distance.
model size / context length 1 2 3 4 5 6 7
160M 0.737 0.606 0.527 0.477 0.445 0.418 0.397
420M 0.753 0.621 0.534 0.480 0.445 0.414 0.391
1.4B 0.768 0.635 0.543 0.484 0.446 0.413 0.388
Table 19: Top-1 accuracy of optimal rule with model scale (Wikipedia, $L^{\infty}$ distance). How top-1 accuracy of the optimal rule varies with model size and rule context length. Optimal rule is selected from $\mathcal{R}^{\textrm{all}}_{7}$ using $L^{\infty}$ distance.
model size / context length 1 2 3 4 5 6 7
160M 25.9 41.2 50.4 55.9 60.4 64.4 67.8
420M 24.7 40.3 50.0 55.7 60.4 64.3 67.7
1.4B 24.3 39.7 49.8 55.5 60.3 64.3 67.7
Table 20: Optimal rule distance with model scale (Wikipedia, $L^{\infty}$ distance). How optimal rule distance varies with model size and rule context length. Optimal rule is selected from $\mathcal{R}^{\textrm{all}}_{7}$ using $L^{\infty}$ distance.
model size context length 1 2 3 4 5 6 7
160M 0.429 0.321 0.253 0.208 0.179 0.154 0.135
420M 0.455 0.346 0.271 0.222 0.190 0.163 0.143
1.4B 0.480 0.369 0.290 0.237 0.202 0.173 0.150
D.4 Rule Selection